A 30 N object is on a 2.0 kg wagon as it is being pulled at a constant velocity by a 10.0 N force 20.0 0 above the positive horizontal. Calculate:

a. The normal force acting on the wagon

b. The frictional force acting on the wagon
1
Expert's answer
2020-04-01T10:18:43-0400

a)


"N=W_O+W_W-F\sin{20}"

"N=30+(2)(9.8)-10\sin{20}=46\ N"

b)


"F_f=F\cos{20}=10\cos{20}=9.4\ N"